OpenStudy (keke):

simplify ; sq root of 11/13

OpenStudy (anonymous):

\[\sqrt{11/13} = (\sqrt{11}/\sqrt{13})*(\sqrt{13}/\sqrt{13}) = \sqrt{143}/13\]